Suppose a random variable x is best described by a normal distribution with μ = 60 and Find the z-score that corresponds to the value x = 63.


1
Expert's answer
2022-03-29T16:00:56-0400

"z=\frac{63-60}{\sigma}=\frac{3}{\sigma}" .